# VegasAfterglow: A High-Performance Framework for Gamma-Ray Burst Afterglows

### v1

@@ Abstract
Gamma-ray bursts (GRBs) are the most luminous astrophysical transients, known to be associated with core collapse of massive stars or mergers of two compact objects such as two neutron stars. They are followed by multi-wavelength afterglow emission originating from the deceleration of the relativistic jets by the ambient medium. The study of after emission offers crucial insights into the physics of relativistic shocks, the properties of the circumburst environment, as well as the physical and geometrical structure of relativistic jets, as well as the viewing geometry of the observer. We present VegasAfterglow, a newly developed, high-performance C++ framework designed for modeling GRB afterglows with flexibility and computational efficiency as keynotes of design. The framework self-consistently solves forward and reverse shock dynamics and calculates synchrotron (including self-absorption or all spectral regimes) and inverse Compton radiation (including Klein–Nishina corrections); it can handle arbitrary user-defined ambient density profiles, central engine activity histories, viewing angles, and the jet structures of energy, Lorentz factor, and magnetization profiles. It supports both relativistic and non-relativistic regimes and includes lateral jet spreading effects. In this paper, we describe the numerical implementation of the framework and assess its computational performance. Our results demonstrate that VegasAfterglow is well-suited for interpreting current and future multi-wavelength observations in the era of multi-messenger astronomy.
@@ 1 Introduction
While early GRB modeling often assumed a uniform “top-hat” jet, with constant energy and Lorentz factor within a narrow opening angle and a sharp cutoff beyond (Rhoads, 1997; Sari et al., 1999), this model only accurately describes on-axis observations. At cosmological distances, selection effects favor such detections because the prompt emission is highly beamed. In this scenario, the afterglow lightcurve (in optical and higher energy bands) typically begins bright and decays monotonically. However, this interpretation breaks down when considering observers located outside the jet core. More realistic models allow for jets with angular structure, where energy and Lorentz factor decrease smoothly with angle from the jet axis (Mészáros et al., 1998; Dai and Gou, 2001; Rossi et al., 2002a; Zhang and Mészáros, 2002a; Kumar and Granot, 2003; Granot and Kumar, 2003; Zhang et al., 2004). In both scenarios, relativistic beaming makes off-axis afterglows observable, typically exhibiting a delayed rise to peak brightness followed by a decay, though the exact light curve shape and timescale vary depending on the prescribed jet structure.
@@
Over the past two decades, numerous numerical and semi-analytical codes have been developed to model GRB afterglows (e.g. Panaitescu et al., 1998; Dermer et al., 2000; Huang et al., 2000; Zhang and Mészáros, 2001b; Granot and Sari, 2002; Granot and Kumar, 2003; van Eerten et al., 2012; Granot et al., 2018; Ryan et al., 2020; Wang et al., 2024). They range from simplified top-hat models to sophisticated hydrodynamical simulations. However, many existing public frameworks lack the flexibility needed to model multi-component or user-defined jet structures, reverse shock emission, synchrotron self-absorption, and inverse Compton (IC) contributions in both the Thomson and Klein-Nishina regimes. In general, analytic and semi-analytic tools are efficient and flexible, allowing for fast light curve generation, but often rely on assumptions like conical geometry, negligible lateral spreading, or approximate blastwave dynamics. Full relativistic hydrodynamic simulations capture complex features such as lateral expansion, shock structure, and jet–medium interaction, but are computationally expensive and impractical for broad parameter inference (e.g., via MCMC sampling).
@@ 2 Jets and Medium
VegasAfterglow models the jet dynamics on a 3D grid of spherical coordinates $(t,\theta,\phi)$, where $t$ is the source frame time, $\theta$ is the polar angle, and $\phi$ is the azimuthal angle. The jet is modeled as a relativistic outflow with a Lorentz factor profile $\Gamma_{0}(\phi,\theta)$, energy profile $\frac{dE}{d\Omega}(\phi,\theta)$, and magnetization profile $\sigma(\phi,\theta)$. In our afterglow calculation, the jet is decomposed into a grid of elements on $\theta$ and $\phi$, and each jet element evolves independently.
@@ 2.1 Jet profiles
[…]
@@
    The Gaussian jet profile (Zhang and Mészáros, 2002a; Kumar and Granot, 2003; Zhang et al., 2004) is defined as
@@
    The power-law jet profile (Mészáros et al., 1998; Dai and Gou, 2001; Rossi et al., 2002b; Zhang and Mészáros, 2002a; Granot and Kumar, 2003) is defined as
    $\Gamma_{0}(\theta)$ $=$ $(\Gamma_{\rm 0}(0)-1)\bigg{(}1+\frac{\theta}{\theta_{\rm c}}\bigg{)}^{-k}+1\,,$ (7)
    $\frac{dE}{d\Omega}(\theta)$ $=$ $\frac{E_{\rm iso}(0)}{4\pi}\bigg{(}1+\frac{\theta}{\theta_{\rm c}}\bigg{)}^{-k}\,,$ (8)
@@ 2.3 Medium
$\rho(\phi,\theta,r)=\frac{A}{r^{k}}\,,$ (13)
where $r_{\rm c}$ and $n_{\rm c}$ are the characteristic radius and corresponding number density of the wind and $k$ is the wind profile index, which is usually adopted as 2 for a constant-velocity wind due to mass flow conservation.
@@ 3.2.1 Reverse Shock Crossing via pressure balance
Combine Eqs. 34 and 38 with this discontinuity condition to get
@@
If the pressure in region 4 (the unshocked jet) is too strong, the reverse shock cannot be formed. Using Eq. 41, one gets the reverse shock generation condition $e_{3}>0$, which is
@@
$\lesssim$ $\frac{8}{3}(\Gamma_{3}^{2}-1)\frac{n_{1}}{n_{4}}\sim\frac{8}{3}(\Gamma_{4}^{2}-1)\frac{n_{1}}{n_{4}}\,,$ (44)
@@ 3.2.2 Reverse Shock Crossing via energy conservation
Eq. 41 assumes that the Lorentz factor and pressure in regions 2 and 3 are constant (thus the Lorentz factor and pressure immediately behind the shock front in the downstream equal the pressure at the discontinuity). This approximation is valid when the reverse shock crossing time is short and the resulting blast wave is thin. However, for a long-duration reverse shock that produces a thick blast wave, this assumption is invalid and the pressure-balance method can lead to significant violations of energy conservation.
@@
Rather than adopting existing mechanical models, we employ an energy conservation method to solve the shock dynamics during the reverse shock crossing phase (Zhang and Mészáros, 2001a, 2002b). The total energy of the system as the blast wave propagates to $r$ can be expressed as the initial energy of the jet plus the rest mass energy of the swept-up medium,
$E_{\rm jet}+E_{\rm medium}=N_{4,0}\Gamma_{4}m_{p}c^{2}(1+\sigma_{4})+N_{2}m_{p}c^{2}\,,$ (45)
where $N_{4,0}$ and $N_{2}=m(r)/m_{p}$ are the proton numbers of the entire jet and region 2, respectively. The total energy in regions 2, 3 and 4 is
$E_{2}+E_{3}+E_{4}$ $=$ $N_{2}[\Gamma_{2}+\Gamma_{\rm eff,2}(\Gamma_{21}-1)]m_{p}c^{2}$ (46)
$+$ $N_{3}[\Gamma_{3}+\Gamma_{\rm eff,3}(\Gamma_{43}-1)]m_{p}c^{2}(1+\sigma_{4})$
$+$ $N_{4}\Gamma_{4}m_{p}c^{2}(1+\sigma_{4})\,,$
@@
$\Gamma_{\rm eff,2}$ $=$ $\frac{\hat{\gamma}_{21}\Gamma_{2}^{2}-\hat{\gamma}_{21}+1}{\Gamma_{2}}\,,$ (47)
$\Gamma_{\rm eff,3}$ $=$ $\frac{\hat{\gamma}_{43}\Gamma_{3}^{2}-\hat{\gamma}_{43}+1}{\Gamma_{3}}\,,$ (48)
are effective Lorentz factors that account for both internal energy and pressure in the energy term of the stress-energy tensor. Similar to previous mechanical models, Equation 46 assumes a constant Lorentz factor profile, $\Gamma$, within the blast wave. Under this assumption, the total internal energy can be expressed as $U=(\Gamma_{\rm ud}-1)Nm_{p}c^{2}$. By applying energy and particle number conservation, where
$E_{\rm jet}+E_{\rm medium}$ $=$ $E_{2}+E_{3}+E_{4}\,,$ (49)
$N_{4,0}$ $=$ $N_{3}+N_{4}\,,$ (50)
and assuming $\Gamma_{3}=\Gamma_{2}=\Gamma_{21}$, we obtain
00 0 $=$ $N_{2}[\Gamma_{3}-1+\Gamma_{\rm eff,2}(\Gamma_{3}-1)]$ (51)
$+$ $N_{3}(1+\sigma_{4})[\Gamma_{3}-\Gamma_{4}+\Gamma_{\rm eff,3}(\Gamma_{43}-1)]\,.$
Here, $N_{2}$, $\Gamma_{4}$, and $\sigma_{4}$ are known quantities, and $N_{3}$ can be expressed as
[…]
$\frac{d\Delta_{3}^{\prime}}{dr}$ $=$ $\Gamma_{3}\frac{\beta_{4}-\beta_{3}}{\beta_{3}}\bigg{(}\frac{\Gamma_{3}n_{3}}{\Gamma_{4}n_{4}}-1\bigg{)}^{-1}\,,$ (53)
@@
where $\Delta_{3}^{\prime}$ is the co-moving width of region 3. With these relations, one can solve $\Gamma_{3}(r)$ from Equation 51 until the reverse shock has fully propagated through the ejecta, which occurs when
$N_{3}(r_{\times})=N_{4}\equiv\frac{E_{\rm jet}}{\Gamma_{4}m_{p}c^{2}(1+\sigma_{4})}\,,$ (54)
@@
where $l$ is the Sedov length and $\Delta=cT$, with $T$ being the duration of the burst. This approach effectively functions as a simplified mechanical model: rather than explicitly integrating the pressure, enthalpy, and density profiles within the blast wave, it relies on the integrated quantities $N_{2}$ and $N_{3}$, which implicitly encode this information and are independent of the internal density structure. This formulation greatly simplifies the traditional mechanical model while retaining its essential physics.
@@
Figure 3 shows examples of the forward and reverse shock dynamics for different shell thicknesses and magnetizations. The upper panels display the evolution of the four-velocities of the forward shock ($\Gamma_{12}$), reverse shock ($\Gamma_{43}$), and blast wave tail ($\Gamma_{3}$) for the unmagnetized case in both the thick-shell regime ($\xi_{0}\ll 1$) and thin-shell regime ($\xi_{0}\gg 1$).
We remark that in the thin shell regime, where $\xi_{0}\gg 1$, the reverse shock remains non-relativistic throughout the time it crosses the entire shell, with a peak Lorentz factor $\Gamma_{43}\sim 1.1$. This is in contrast to the mildly relativistic value $\Gamma_{43}\sim 7/4$ derived in Sari and Piran (1995); Kobayashi (2000). Several factors contribute to this discrepancy:
@@
    Most importantly, as the reverse shock (characterized by $\Gamma_{43}$) strengthens, the bulk Lorentz factor $\Gamma_{3}$ decreases toward $\sim\Gamma_{4}/2$ during the crossing. This evolution gradually changes the scaling of $\Gamma_{43}-1\propto t^{3}$ (typical for the thin shell regime) to $\Gamma_{43}-1\propto t^{1/4}$ (typical for the thick shell regime), thereby slowing the growth of $\Gamma_{43}$ in the later stages of the crossing (see upper right panel of Figure 3).
@@
    The energy conservation method adopted in this work obtains a slightly lower $\Gamma_{43}$ than the traditional pressure balance method.
@@
Figure 4: Numerical results for the reverse shock crossing time, $t_{\times}$, as a function of upstream magnetization, $\sigma$, for different shell thicknesses, based on Eqs. 52 and 54. The green dashed lines indicate the analytical scaling derived in Zhang and Kobayashi (2005), while the orange horizontal dashed lines represent the engine duration.
@@
Figure 4 shows our numerical results for the reverse shock crossing time, $t_{\times}$, as a function of upstream magnetization, $\sigma$, for different shell thicknesses, obtained by solving Equation 52 and 54. We remark that $t_{\times}$ does not strictly follow the analytical scaling derived in Zhang and Kobayashi (2005), where the deceleration radius coefficient was estimated as $C_{\gamma}\sim(1+\sigma)^{-1/3}$ and the shock crossing radius coefficient as $C_{\Delta}\sim(1+\sigma)^{-1/2}$. Under this assumption, the shock crossing time scales as $t_{\times}\sim t_{\gamma}\left(\frac{C_{\Delta}}{C_{\gamma}}\right)^{4}\propto t_{\gamma}/(1+\sigma)^{2/3}$, where $t_{\gamma}\sim l/(2\Gamma_{4}^{8/3}c)$ is the deceleration time in the thin-shell regime for an unmagnetized upstream with Sedov length $l$.
@@
The original derivation assumes that $C_{\Delta}=\left(1-\frac{\Gamma_{4}n_{4}}{\Gamma_{3}n_{3}}\right)^{1/2}$ is effectively independent of $\Gamma_{43}$ and scales as $(1+\sigma)^{-1/2}$. This assumption holds for mildly and ultra-relativistic reverse shocks, as shown in Figure 3 of Zhang and Kobayashi (2005). However, based on our fully analytical shock jump conditions (Eqs. 14, 15, and 26), we find that $C_{\Delta}$ deviates from the $(1+\sigma)^{-1/2}$ scaling when $\Gamma_{43}-1$ is small, which is typically the case during the early stage in thin-shell regime.
@@
Figure 5: The reverse shock crossing radius coefficient $C_{\Delta}$ (See definition in Eq. 39 of Zhang and Kobayashi (2005)) as a function of upstream magnetization, $\sigma$, and reverse shock strength, $\Gamma_{34}-1$. The upstream Lorentz factor is set to 2000. The red dashed line shows the analytical scaling derived in Zhang and Kobayashi (2005).

### v3

@@ Abstract
Gamma-ray bursts (GRBs) are the most luminous astrophysical transients, known to be associated with core collapse of massive stars or mergers of two compact objects such as two neutron stars. They are followed by multi-wavelength afterglow emission originating from the deceleration of the relativistic jets by the ambient medium. The study of afterglow emission offers crucial insights into the physics of relativistic shocks, the properties of the circumburst environment, the physical and geometrical structure of relativistic jets, as well as the viewing geometry of the observer. We present VegasAfterglow, a newly developed, high-performance C++ framework designed for modeling GRB afterglows with flexibility and computational efficiency as key features of design. The framework self-consistently solves forward and reverse shock dynamics and calculates synchrotron (including self-absorption or all spectral regimes) and inverse Compton radiation (including Klein–Nishina corrections); it can handle arbitrary user-defined ambient density profiles, central engine activity histories, viewing angles, and the jet structures of energy, Lorentz factor, and magnetization profiles. It supports both relativistic and non-relativistic regimes and includes lateral jet spreading effects. In this paper, we describe the numerical implementation of the framework and assess its computational performance. Our results demonstrate that VegasAfterglow is well-suited for interpreting current and future multi-wavelength observations in the era of multi-messenger astronomy.
@@ 1 Introduction
While early GRB modeling often assumed a uniform “top-hat” jet, with constant energy and Lorentz factor within a narrow opening angle and a sharp cutoff beyond (Rhoads, 1997; Sari et al., 1999), this model only accurately describes on-axis observations. At cosmological distances, selection effects favor such detections because the prompt emission is highly beamed. In this scenario, the afterglow light curve (in optical and higher energy bands) typically begins bright and decays monotonically. However, this interpretation breaks down when considering observers located outside the jet core. More realistic models allow for jets with angular structure, where energy and Lorentz factor decrease smoothly with angle from the jet axis (Mészáros et al., 1998; Dai and Gou, 2001; Rossi et al., 2002a; Zhang and Mészáros, 2002; Kumar and Granot, 2003; Granot and Kumar, 2003; Zhang et al., 2004). In both scenarios, relativistic beaming makes off-axis afterglows observable, typically exhibiting a delayed rise to peak brightness followed by a decay, though the exact light curve shape and timescale vary depending on the prescribed jet structure.
@@
Over the past two decades, numerous numerical and semi-analytical codes have been developed to model GRB afterglows (e.g. Panaitescu et al., 1998; Dermer et al., 2000; Huang et al., 2000; Zhang and Mészáros, 2001b; Granot and Sari, 2002; Granot and Kumar, 2003; van Eerten et al., 2012; Lei et al., 2016; Granot et al., 2018; Ryan et al., 2020; Wang et al., 2024; Nedora et al., 2024; Kusafuka et al., 2025). They range from simplified top-hat models to sophisticated hydrodynamical simulations. However, many existing public frameworks lack the flexibility needed to model multi-component or user-defined jet structures, reverse shock emission, synchrotron self-absorption, and inverse Compton (IC) contributions in both the Thomson and Klein-Nishina regimes. In general, analytic and semi-analytic tools are efficient and flexible, allowing for fast light curve generation, but often rely on assumptions like conical geometry, negligible lateral spreading, or approximate blast wave dynamics. Full relativistic hydrodynamic simulations capture complex features such as lateral expansion, shock structure, and jet–medium interaction, but are computationally expensive and impractical for broad parameter inference (e.g., via MCMC sampling).
@@ 2 Jets and Medium
VegasAfterglow models jet dynamics on a 3D grid of spherical coordinates $(t,\theta,\phi)$, where $t$ is the source frame time, $\theta$ is the polar angle, and $\phi$ is the azimuthal angle. The jet is modeled as a relativistic outflow with a Lorentz factor profile $\Gamma_{0}(\phi,\theta)$, energy profile $\frac{dE}{d\Omega}(\phi,\theta)$, and magnetization profile $\sigma(\phi,\theta)$. In our afterglow calculation, the jet is decomposed into a grid of elements on $\theta$ and $\phi$, and each jet element evolves independently.
@@ 2.1 Jet profiles
    $\Gamma_{0}(\theta)$ $=$ $\left\{\begin{aligned} &\Gamma_{\rm 0},\quad\theta\leq\theta_{\rm c}\\ &1,\quad\theta>\theta_{\rm c}\end{aligned}\right$.\,, (1)
    $\frac{dE}{d\Omega}(\theta)$ $=$ $\left\{\begin{aligned} &\frac{E_{\rm iso}}{4\pi},\quad\theta\leq\theta_{\rm c}\\ &0,\quad\theta>\theta_{\rm c}\end{aligned}\right$.\,, (2)
    $\sigma(\theta)$ $=$ $\left\{\begin{aligned} &\sigma_{\rm 0},\quad\theta\leq\theta_{\rm c}\\ &0,\quad\theta>\theta_{\rm c}\end{aligned}\right.\,$. (3)
@@
    The Gaussian jet profile (Zhang and Mészáros, 2002; Kumar and Granot, 2003; Zhang et al., 2004) is defined as
@@
    The power-law jet profile (Mészáros et al., 1998; Dai and Gou, 2001; Rossi et al., 2002b; Zhang and Mészáros, 2002; Granot and Kumar, 2003) is defined as
    $\Gamma_{0}(\theta)$ $=$ $(\Gamma_{\rm 0}(0)-1)\bigg(1+\bigg(\frac{\theta}{\theta_{\rm c}}\bigg)^{k}\bigg)^{-1}+1\,$, (7)
    $\frac{dE}{d\Omega}(\theta)$ $=$ $\frac{E_{\rm iso}(0)}{4\pi}\bigg(1+\bigg(\frac{\theta}{\theta_{\rm c}}\bigg)^{k}\bigg)^{-1}\,$, (8)
@@ 2.3 Medium
$\rho(\phi,\theta,r)=\rho_{c}\bigg(\frac{r}{r_{c}}\bigg)^{-k}\,$, (13)
where $r_{\rm c}$ and $\rho_{\rm c}$ are the characteristic radius and corresponding density of the wind and $k$ is the wind profile index, which is usually adopted as 2 for a constant-velocity wind due to mass flow conservation.
@@ 3.2.1 Reverse Shock Crossing via pressure balance
Combine Equation 34 and 38 with this discontinuity condition to get
@@
If the pressure in region 4 (the unshocked jet) is too strong, the reverse shock cannot be formed. Using Equation 41, one gets the reverse shock generation condition $e_{3}>0$, which is
@@
$\lesssim$ $\frac{8}{3}(\Gamma_{3}^{2}-1)\frac{n_{1}}{n_{4}}\sim\frac{8}{3}(\Gamma_{4}^{2}-1)\frac{n_{1}}{n_{4}}\,$,
@@ 3.2.2 Reverse Shock Crossing via energy conservation
Equation 41 assumes that the Lorentz factor and pressure in regions 2 and 3 are constant (thus the Lorentz factor and pressure immediately behind the shock front in the downstream equal the pressure at the discontinuity). This approximation is valid when the reverse shock crossing time is short and the resulting blast wave is thin. However, for a long-duration reverse shock that produces a thick blast wave, this assumption is invalid and the pressure-balance method can lead to significant violations of energy conservation.
@@
Rather than adopting existing mechanical models, we employ an energy conservation method to solve the shock dynamics during the reverse shock crossing phase. During the shock crossing phase, the total energy in the blast wave, including region 2 and region 3, can be expressed as
$\Gamma_{2}m_{2}c^{2}+\Gamma_{\rm eff,2}U_{2}+\Gamma_{3}m_{3}c^{2}+\Gamma_{\rm eff,2}U_{2}$ (44)
@@
$\Gamma_{\rm eff,2}$ $=$ $\frac{\hat{\gamma}_{2}\Gamma_{2}^{2}-\hat{\gamma}_{2}+1}{\Gamma_{2}}\,$, (45)
$\Gamma_{\rm eff,3}$ $=$ $\frac{\hat{\gamma}_{34}\Gamma_{3}^{2}-\hat{\gamma}_{34}+1}{\Gamma_{3}}\,.$ (46)
During the shock crossing phase, the forward shock propagates into the interstellar medium and compresses the material into region 2, while the reverse shock compresses the material from region 4 into region 3. Therefore, the blast wave dynamics equation can be written as
$d[\Gamma_{2}m_{2}c^{2}+\Gamma_{\rm eff,2}U_{2}+\Gamma_{3}m_{3}c^{2}+\Gamma_{\rm eff,2}U_{2}]=c^{2}dm_{2}$
$+c^{2}\Gamma_{4}dm_{3}+\Gamma_{\rm eff,2}dU_{\rm rad,2}+\Gamma_{\rm eff,3}dU_{\rm rad,3}+L_{\rm inj}dt$ (47)
where $dU_{\rm rad,2}$ and $dU_{\rm rad,3}$ are the internal energy losses due to radiation in regions 2 and 3, respectively, and $L_{\rm inj}$ is the external energy injection luminosity. The internal energy evolution is
$dU_{\rm i}=dU_{\rm sh,i}+dU_{\rm ad,i}+dU_{\rm rad,i}$ (48)
which includes shock heating $dU_{\rm sh,i}$, adiabatic cooling $dU_{\rm ad,i}$, and radiation $dU_{\rm rad,i}$, where
$dU_{\rm sh,2}$ $=$ $(\Gamma_{2}-1)c^{2}dm_{2}$ (49)
$dU_{\rm rad,2}$ $=$ $-\epsilon_{2}dU_{\rm sh,2}=-\epsilon_{\rm rad,2}\epsilon_{e,2}dU_{\rm sh,2}$ (50)
$dU_{\rm ad,2}$ $=$ $-pdV_{2}^{\prime}=-(\hat{\gamma}_{2}-1)\frac{U_{2}}{V_{2}^{\prime}}dV_{2}^{\prime}$ (51)
$=$ $-(\hat{\gamma}_{2}-1)U_{2}\bigg(2\frac{dr}{r}+\frac{d\Delta_{2}^{\prime}}{\Delta_{2}^{\prime}}\bigg)$
$dU_{\rm sh,3}$ $=$ $(\Gamma_{34}-1)c^{2}dm_{3}$ (52)
$dU_{\rm rad,3}$ $=$ $-\epsilon_{3}dU_{\rm sh,3}=-\epsilon_{\rm rad,3}\epsilon_{e,3}dU_{\rm sh,3}$ (53)
$dU_{\rm ad,3}$ $=$ $-pdV_{3}^{\prime}=-(\hat{\gamma}_{34}-1)\frac{U_{3}}{V_{3}^{\prime}}dV_{3}^{\prime}$ (54)
$=$ $-(\hat{\gamma}_{34}-1)U_{3}\bigg(2\frac{dr}{r}+\frac{d\Delta_{3}^{\prime}}{\Delta_{3}^{\prime}}\bigg)$
where $V_{i}^{\prime}=4\pi r^{2}\Delta_{i}^{\prime}$ is the comoving volume in region ii [^3] For adiabatic cooling, we adopt a more accurate description of shell evolution instead of the commonly used $r/\Gamma$ approximation in the literature., $\Delta^{\prime}_{i}$ is the comoving shell thickness in region $i$, and the radiation efficiency in region $i$ can be expressed as
$\epsilon_{\rm rad,i}=\left\{\begin{aligned} &1,\quad&\gamma_{c,i}<\gamma_{m,i}&\\ &\bigg(\frac{\gamma_{c,i}}{\gamma_{m,i}}\bigg)^{2-p},\quad&\gamma_{c,i}>\gamma_{m,i}&,\quad p_{i}>2&\end{aligned}\right$.\,, (55)
(see the definitions of $\gamma_{m}$, $\gamma_{c}$ and $\epsilon_{e}$ in Section 4).
To close these equations, we also need the shell evolution equation:
$\frac{d\Delta_{2}^{\prime}}{dt}$ $=$ $c_{s,2}\frac{dt^{\prime}}{dt}$ (56)
$\frac{d\Delta_{3}^{\prime}}{dt}$ $=$ $\Gamma_{3}\frac{\beta_{4}-\beta_{3}}{\beta_{3}}\bigg(\frac{\Gamma_{3}n_{3}}{\Gamma_{4}n_{4}}-1\bigg)^{-1}\frac{dr}{dt}\,$, (57)
where
$c_{s,2}=c\sqrt{\frac{\hat{\gamma}_{2}p}{\rho_{0}c^{2}+\frac{\hat{\gamma}_{2}}{\hat{\gamma}_{2}-1}p}}=c\sqrt{\frac{\hat{\gamma}_{2}(\hat{\gamma}_{2}-1)(\Gamma_{2}-1)}{1+\hat{\gamma}_{2}(\Gamma_{2}-1)}}$ (58)
is the sound speed in region 2, and the mass accumulation equations are
$\frac{dm_{2}}{dt}$ $=$ $4\pi r^{2}\rho_{1}\frac{dr}{dt}$ (59)
$\frac{dm_{3}}{dt}$ $=$ $4\pi r^{2}n_{3}m_{p}\frac{d\Delta_{3}^{\prime}}{dt}$ (60)
By assuming $\Gamma_{2}=\Gamma_{3}$, ${dr_{2}}/{dt}={dr_{3}}/{dt}={c\beta_{2}}/(1-\beta_{2})$ during the shock crossing phase, and putting everything together, one obtains
$\frac{d\Gamma_{2}}{dt}$ $=$ $-\frac{(\Gamma_{2}-1)c^{2}\frac{dm_{2}}{dt}+\Gamma_{\rm eff,2}(\frac{dU_{\rm sh,2}}{dt}+\frac{dU_{\rm ad,2}}{dt})}{(m_{2}+m_{3})c^{2}+\frac{d\Gamma_{\rm eff,2}}{d\Gamma_{2}}U_{2}+\frac{d\Gamma_{\rm eff,3}}{d\Gamma_{2}}U_{3}}$ (61)
- $\frac{(\Gamma_{2}-\Gamma_{4})c^{2}\frac{dm_{3}}{dt}+\Gamma_{\rm eff,3}(\frac{dU_{\rm sh,3}}{dt}+\frac{dU_{\rm ad,3}}{dt})}{(m_{2}+m_{3})c^{2}+\frac{d\Gamma_{\rm eff,2}}{d\Gamma_{2}}U_{2}+\frac{d\Gamma_{\rm eff,3}}{d\Gamma_{2}}U_{3}}$
$+$ $\frac{L_{\rm inj}}{(m_{2}+m_{3})c^{2}+\frac{d\Gamma_{\rm eff,2}}{d\Gamma_{2}}U_{2}+\frac{d\Gamma_{\rm eff,3}}{d\Gamma_{2}}U_{3}}$
$\frac{dU_{2}}{dt}$ $=$ $(1-\epsilon_{2})(\Gamma_{2}-1)c^{2}\frac{dm_{2}}{dt}$ (62)
- $(\hat{\gamma}_{2}-1)U_{2}\bigg(\frac{2}{r}\frac{dr}{dt}+\frac{1}{\Delta_{2}^{\prime}}\frac{d\Delta_{2}^{\prime}}{dt}\bigg)$
$\frac{dU_{3}}{dt}$ $=$ $(1-\epsilon_{3})(\Gamma_{34}-1)c^{2}\frac{dm_{3}}{dt}$ (63)
- $(\hat{\gamma}_{34}-1)U_{3}\bigg(\frac{2}{r}\frac{dr}{dt}+\frac{1}{\Delta_{3}^{\prime}}\frac{d\Delta_{3}^{\prime}}{dt}\bigg)$
@@
Equation 61 describes the blast wave dynamics until the reverse shock has fully propagated through the ejecta, which occurs when
$m_{3}(r_{\times})=m_{4}\equiv\frac{E_{\rm jet}}{\Gamma_{4}c^{2}(1+\sigma_{4})}\,$, (64)
@@
where $l$ is the Sedov length and $\Delta=cT$, with $T$ being the duration of the burst. This approach effectively functions as a simplified mechanical model: rather than explicitly integrating the pressure, enthalpy, and density profiles within the blast wave, it relies on the total energy in region 2 and region 3. This formulation greatly simplifies the traditional mechanical model while retaining its essential physics.
@@
Figure 3 shows examples of the forward and reverse shock four-velocity evolution for different shell thicknesses and magnetizations. The upper panels display the evolution of the four-velocities of the forward shock ($\Gamma_{12}$), reverse shock ($\Gamma_{43}$), and blast wave tail ($\Gamma_{3}$) for the unmagnetized case in both the thick-shell regime ($\xi_{0}\ll 1$) and thin-shell regime ($\xi_{0}\gg 1$).
We remark that in the thin shell regime, where $\xi_{0}\gg 1$, the reverse shock remains non-relativistic throughout the time it crosses the entire shell, with a peak Lorentz factor $\Gamma_{43}\sim 1.02$ ($\Gamma_{34}\beta_{34}\sim 0.2$). This is in contrast to the mildly relativistic value $\Gamma_{43}\sim 7/4$ derived in Sari and Piran (1995) and Kobayashi (2000). Several factors contribute to this discrepancy:
@@
    As the reverse shock (characterized by $\Gamma_{43}$) strengthens, the bulk Lorentz factor $\Gamma_{3}$ decreases toward $\sim\Gamma_{4}/2$ during the crossing. This evolution gradually changes the scaling of $\Gamma_{43}-1\propto t^{3}$ (typical for the thin shell regime) to $\Gamma_{43}-1\propto t^{1/4}$ (typical for the thick shell regime), thereby slowing the growth of $\Gamma_{43}$ in the later stages of the crossing (see upper right panel of Figure 3).
@@
    The energy-conservation method adopted in this work yields a lower $\Gamma_{43}$ than the traditional pressure-balance method. This difference arises because the commonly used internal energy approximation, $U_{3}=(\Gamma_{34}-1)m_{3}c^{2}$, overestimates the actual internal energy in region 3. The exact internal energy is given by the integral form of Equation 63, in which the early shock heating is weaker. Moreover, adiabatic cooling further reduces the internal energy relative to $(\Gamma_{34}-1)m_{3}c^{2}$. A smaller $U_{3}$ in the total energy budget of Equation 44 leads to a higher $\Gamma_{3}$ during the reverse shock crossing, and consequently, the reverse shock is significantly weakened.
@@
Figure 4: Numerical results for the reverse shock crossing time, $t_{\times}$, as a function of upstream magnetization, $\sigma$, for different shell thicknesses, based on Equation 60 and 64. The green dashed lines indicate the analytical scaling derived in Zhang and Kobayashi (2005), while the orange horizontal dashed lines represent the engine duration.
@@
Figure 4 shows our numerical results for the reverse shock crossing time, $t_{\times}$, as a function of upstream magnetization, $\sigma$, for different shell thicknesses, obtained by solving Equation 60 and 64. We remark that $t_{\times}$ does not strictly follow the analytical scaling derived in Zhang and Kobayashi (2005), where the deceleration radius coefficient was estimated as $C_{\gamma}\sim(1+\sigma)^{-1/3}$ and the shock crossing radius coefficient as $C_{\Delta}\sim(1+\sigma)^{-1/2}$. Under this assumption, the shock crossing time scales as $t_{\times}\sim t_{\gamma}\left(\frac{C_{\Delta}}{C_{\gamma}}\right)^{4}\propto t_{\gamma}/(1+\sigma)^{2/3}$, where $t_{\gamma}\sim l/(2\Gamma_{4}^{8/3}c)$ is the deceleration time in the thin-shell regime for an unmagnetized upstream with Sedov length $l$.
@@
The original derivation assumes that $C_{\Delta}=\left(1-\frac{\Gamma_{4}n_{4}}{\Gamma_{3}n_{3}}\right)^{1/2}$ is effectively independent of $\Gamma_{43}$ and scales as $(1+\sigma)^{-1/2}$. This assumption holds for mildly and ultra-relativistic reverse shocks, as shown in Figure 3 of Zhang and Kobayashi (2005). However, based on our fully analytical shock jump conditions (Equation 14, 15, and 26), we find that $C_{\Delta}$ deviates from the $(1+\sigma)^{-1/2}$ scaling when $\Gamma_{43}-1$ is small, which is typically the case during the early stage in thin-shell regime.
@@
Figure 5: The reverse shock crossing radius coefficient $C_{\Delta}$ (See definition in Equation 39 of Zhang and Kobayashi (2005)) as a function of upstream magnetization, $\sigma$, and reverse shock strength, $\Gamma_{34}-1$. The upstream Lorentz factor is set to 2000. The red dashed line shows the analytical scaling derived in Zhang and Kobayashi (2005).